AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
AIDA
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 15:07:31 2025
Author
*
:
Subject
*
:
<p>19/02/15</p> <p>1000: Check system gives same resolution as yesterday, removing only the Cu braiding connecting and grounding all adapter PCBs (also moved bias braid back to nnaida3).</p> <p> Pulser FWHM in each FEE... nnaida1 ~800ch, nnaida3 ~110ch, nnaida6 ~2300ch, nnaida8 ~950ch. Rate in 6 and 8 ~5kHz per strip.</p> <p> Re-added Cu braids, clipping on to LEMO connectors of adapter PCBs.</p> <p> FWHM and rate/strip in nnaida1,3,6,8 now (respectively): ~600ch and 100Hz, ~120ch and 25Hz, ~2100ch and 6kHz, 1000ch and 6kHz. => small reduction in nnaida1+6</p> <p> nnaida8 also has non-Gaussian pulser peaks.</p> <p> No clear effect, but minor improvement and no negative effects mean we will keep Cu braids attached.</p> <p>1045: Grounded nose to adapter PCB attached to nnaida7+8. No clear effect.</p> <p>1055: Changed pre-amp reference settings in nnaida1+8.</p> <p> nnaida1 0x30 -> 0x60. nnaida8 0xb2 -> 0x82.</p> <p> No clear change. Reverted back to original settings.</p> <p>1130: Changed to detector BB18 2977-20 and added 241Am (1uCi) and 207Bi (few uCi) sources 2cm from front of detector (see attached photos for arrangement).</p> <p> Bias set at -100V with leakage current -10.9uA.</p> <p> Pulser peak FWHMs for each nnaida are: #1 ~600ch, #3 ~150ch, #6 ~1300ch, #8 ~ 700ch. Minor improvement over previous detector.</p> <p>1345: nnaida3 still much more well behaved than other nnaida. To test if kapton cables are source of issue, swapped kaptons for nnaida3+8 at FEE end.</p> <p> Peak FWHMs and rates for nnaida3+8 having swapped cables are: #3 ~700ch and 300Hz, #8 ~200ch and 30Hz.</p> <p> => noise moved with kapton! Kaptons could be the issue. (For completeness, FWHM in nnaida1 ~800ch and nnaida6 ~1100ch.)</p> <p>1435: Disconnected kaptons at detector end (still connected to FEE adapters).</p> <p> Pulser FWHMs and rates/strip per strip now: nnaida1 - ~250ch and 25Hz, nnaida3 - ~120ch and 25Hz, nnaida6 - ~160ch and 25Hz, nnaida8 - ~120ch and 25Hz.</p> <p> Waveforms for all nnaida looked similar and with little noise => no way of telling a 'bad' cable from a 'good' one, just from FEE + cable (without detector).</p> <p> Current cable configuration/pairings: nnaida1-kapton A, 3-B, 6-C, 8-D.</p> <p> Kapton B shows no alpha/beta spectrum from sources in 1.1.L, a problem which tracked across to nnaida8 when cables were swapped => suggesting break in cable for this channel.</p> <p> Waveforms in all nnaida look relatively noise free (nnaida1 slightly noisier) and as expected.</p> <p>1520: Changed some of the kaptons to check all were functioning properly. Kept nnaida3-kapton B combo as this was working as expected from DL tests.</p> <p> Changed nnaida1 A->E (neither isolated), nnaida6 C->F, nnaida8 D->G.</p> <p> After removal. noticed outgoing kaptons C+D did not have "Isolated" label on them (B does).</p> <p> Without connecting new cables to detector, FWHMs and rates/strip in all nnaida ~130-150ch and ~25Hz, respectively.</p> <p>1545: Connected kaptons to detector 2977-20 and with alpha+beta sources 2cm in front of detector.</p> <p> Bias connected via nnaida6 (-ve core) and nnaida3 (low V ref braid). nnaida3+8 grounded.</p> <p> nnaida-kapton pairings now as follows: nnaida1-kapton E, 3-B, 6-F, 8-G.</p> <p> Pulser FWHMs and rates/strip now: nnaida 1 - ~190ch and ~60Hz, nnaida3 - ~250ch and ~60Hz, nnaida6 - ~180ch and ~40Hz, nnaida8 - ~350ch and ~100Hz.</p> <p> Above rates seem sensible considering we have 25Hz pulser going into FEEs, plus alpha and beta sources.</p> <p> MUCH better than previous! (apart from nnaida3 which remains about the same)</p> <p> Further more, rate spectrum for each nnaida shows slow decline in rate with increasing bin number. Expected as 241Am placed centrally and outer strips will see lower rate as more alphas are stopped due to path travelled.</p> <p>1645: Saving data to file with system configuration: Am+Bi sources, EBFG kaptons, Cu braid grounding between adapters and nose grounded to nnaida7+8 adapter.</p> <p> File - /TapeData/Feb2015/Run1_*.gz.</p> <p> Run stopped at 1715.</p> <p>17:40: Providing one pulser signal to BRIKEN VME crate, which is in the same main bias supply as AIDA. Start a new run to check things are not changed </p> <p> File - /TapeData/Feb2015/Run3_*.gz.</p> <p>19.42 File /TapeDate/R3_1.gz closed (985+303Mb)</p> <div> </div>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown